OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle sin(π/2)cos(π/7)+cos(π/2)sin(π/7)

OpenStudy (mathmate):

hint: use the identity sin(A+B)=sinAcosB+cosAsinB

OpenStudy (michele_laino):

please use the formula of addition of sin x, namely: \[\sin(x+y)=\sin x \cos y+\cos x \sin y\]

OpenStudy (anonymous):

Would I then plug in the numbers accordingly to the formula, or have I got the wrong idea?

OpenStudy (michele_laino):

you can use this, for example: \[x=\frac{ \pi }{ 2 }, y=\frac{ \pi }{ 7 }\]

OpenStudy (anonymous):

So like I said, I would plug these numbers you posted into the formula, correct?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

in the left side of my formula

OpenStudy (anonymous):

Ok. Then onthe right side, the x and y values would be inversed, correct?

OpenStudy (michele_laino):

where you see x, then you insert pi/2, and where you see y then insert pi/7

OpenStudy (anonymous):

I wouldn't need to invert the values for the right side? They get plugged in the same way for both sides?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Ok! Thank you so very much!!

OpenStudy (michele_laino):

Thank you!